OpenStudy (anonymous):

Write the following in summation notation:5+8+11+......+71+74

OpenStudy (anonymous):

Summation of i + 3, ( i = 5 -> 74)

OpenStudy (anonymous):

\[\sum_{3}^{72}i+3\]

OpenStudy (anonymous):

i am sorry; the bottom number should be 2

OpenStudy (anonymous):

2+3=5 and 71 (not 72) + 3= 74

OpenStudy (anonymous):

my basic math is terrible!

OpenStudy (anonymous):

why do you use i instead of n? Just asking because usually i is reserved for complex notation.

OpenStudy (anonymous):

@lisapeajac

OpenStudy (turingtest):

this is wrong I think...

OpenStudy (anonymous):

I used i because that is what the problem had in it; sometimes it is just a variable

OpenStudy (turingtest):

\[\sum_{i=0}^{23}5+3i\]is the answer

OpenStudy (anonymous):

Thanks! much better

OpenStudy (anonymous):

I like to use i for "index", its personal preference.

OpenStudy (anonymous):

\[\sum_{n=2}^{25}3n-1\]